OpenStudy (anonymous):

If 3(m – 2) + 4 – m = 18, what is the value of m?

OpenStudy (anonymous):

1) Distribute. 3(m - 2) + 4 - m = 18 3m - 6 + 4 - m = 18 Combine like terms. 2m - 2 = 18 Add 2. 2m = 20 Divde both sides by 2. m = 10

OpenStudy (anonymous):

Distributive property: a(b + c) = ab + ac